The forward rate of a reaction from S --> P is 10^-9 per second and the reverse reaction is 10^-12 per second. Addition of an enzyme enhances this reaction by 100 fold. What will the reverse rate constant be? Show all your work and explanation please.
Answer #1

Keq=rate constant of the forward rection/rate constant of the backward reaction

Rate of a reaction is directly proportional to the rate contant

Rate for forward reaction k1

10-9 k1

Rate for reverse reaction k2

Therefore, 10-12 k2

Therefore, keq (uncatalyzed) 10-9/10-12=1000

Therefore, k1=k2*1000

Now, for catalyzed, k2*100=k1*100/1000

Therefore,k2(catalyzed)=100*10-12=10-10

Keq for catalyzed and uncatalyzed reaction is the same as it is not changed by the use of a catalyst

Thank You!